OpenStudy (anonymous):

find the slope of the line that contains the points (-2,8) and (1,2) A. 1 over 2 B. 1 C. - 1 over 2 D. -2

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

i hope you are right about these :/ or ill flippin keel my self

OpenStudy (anonymous):

Im always right :D

OpenStudy (triciaal):

slope = change in y / corresponding change in x

OpenStudy (anonymous):

so its d?

OpenStudy (triciaal):

change in y =2-8/change in x =1--2 -6/3 = -2